OpenStudy (anonymous):

Which of the following cannot be the sum of the angle measures of a convex polygon: 1,530 degrees, 3,420 degrees, 6,480 degrees, 4,500 degrees,

OpenStudy (cwrw238):

sum of angles = 180( n -2)

OpenStudy (cwrw238):

n must be a whole number so equate this formula , in turn to the 4 numbers you have and calculate n. if n is not a whole number then you have your answer

OpenStudy (cwrw238):

e. g 3420 180(n - 2) = 3420 180n - 360 = 3420 180n = 3780 n = 3780 / 180 = 21 so 3420 can be the sum of a polygon

OpenStudy (cwrw238):

ok?